OpenStudy (anonymous):

what power to 1/3 = -9

OpenStudy (anonymous):

the power is 1/3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Except... it would have to be to the - 1/3 in order to generate a negative number?

OpenStudy (anonymous):

here is the problem (1/3)^x=-9

OpenStudy (anonymous):

The power of (1/3) (ie. x^(1/3)) is the same as taking the cubed root of the number. Taking a negative power results in a fraction (ie. x^-1 = 1/x). So... \[\sqrt[3]{x}=-9\] \[x=(-9)^3=-729\]

OpenStudy (anonymous):

Misinterpreted the problem. Upon further looking, it seems this problem is not easy at all as it includes logs and imaginary numbers. The answer can be found at http://www.wolframalpha.com/input/?i=%281%2F3%29^x%3D-9

OpenStudy (anonymous):

thank you so much!

jhonyy9 (jhonyy9):

1/3 = -9 3*(-1)= (3i)*2 - so from 1/3 on indifferent power you not can making -9 never

OpenStudy (anonymous):

there is no way of raising \[\frac{1}{3}\] to a power and getting a negative number.